1. The current pretax income for Stapler Inc is $67,000 (tax rate is 24%), with an average asset base of $230,000, and an expected return of 15 percent or higher. The ROI for Stapler Inc would amount to:

3. Lynn Corporation expected to use 1.1 direct labor hours to produce one unit of their product, at a rate of $12/DLH. Actual results for last year indicate that they sold 420,000 units, where their direct labor workforce actually worked 500,000 hours at a rate of $13.25/DLH. What is the Direct Labor Rate Variance?
Answer #1
1.
Net income = Pretax income * ( 1 - Tax% ) = 67000 * ( 1 - 24% ) 50920
ROI = Net income / Average asset = 50920 / 230000 22.1%

3.

Direct labor rate variance = ( Actual rate - Standard rate ) * Actual hours = ( 13.25 - 12 ) * 500000 625000 Unfavorable